The population growth rate in 2015 was about ____.
A) 0.74% B) 0.58% C) 0.66% D) 0.62%

Table GDP.Statl: Vital Demographic and Economic Statistics for The Earthly Republic in Trillion Economic in Million Population 320.99 318.86 316.50 314.11 Per Real GDP 16.30 6.29 15.92 15.43 Capita Per Capita Year 2015 2014 2013 2012 Growth Rate (%) 0.06% 2.32% 3.18% 1.58% GDP 50,780 51.088 50.300 49,123 Income -0.60% 1.57% 2.40% 0.81%
Answer #1

Ans is C

The population growth rate in 2015 = (Population in 2015 - Population in 2014) ÷ Population in 2014

Population in 2015 -Population in 2014 = ( 320.99 - 318.86)Million = 2.13 million

Population growth rate = 2.13 million/ Population in 2014

= 2130000÷318860000 = 0.00668

Population growth rate in percentage = 0.00668×100 = 0.668%

So ans is C 0.66%